# Multi-Robot Motion Planning with Diffusion Models

### v1

@@ Abstract
Diffusion models have recently been successfully applied to a wide range of robotics applications for learning complex multi-modal behaviors from data. However, prior works have mostly been confined to single-robot and small-scale environments due to the high sample complexity of learning multi-robot diffusion models. In this paper, we propose a method for generating collision-free multi-robot trajectories that conform to underlying data distributions while using only single-robot data. Our algorithm, Multi-robot Multi-model planning Diffusion (MMD), does so by combining learned diffusion models with classical search-based techniques—generating data-driven motions under collision constraints. Scaling further, we show how to compose multiple diffusion models to plan in large environments where a single diffusion model fails to generalize well. We demonstrate the effectiveness of our approach in planning for dozens of robots in a variety of simulated scenarios motivated by logistics environments. View video demonstrations in our supplementary material, and our code at: https://github.com/yoraish/mmd.
@@ 1 Introduction
In this paper, we propose a data-efficient and scalable multi-robot diffusion planning algorithm, Multi-robot Multi-model planning Diffusion (MMD), that addresses both these challenges by combining constraint-based MAPF planners with diffusion models. Importantly, our approach calls for learning only single-robot diffusion models, which does away with the difficulty of obtaining multi-robot interaction data and breaks the curse of dimensionality. MMD generates collision-free trajectories by constraining single-robot diffusion models using our novel spatio-temporal guiding functions and choosing constraint placement via strategies inspired by MAPF algorithms. Our contributions in this paper are threefold:
  • We propose a data-efficient framework for multi-robot diffusion planning inspired by constraint-based search algorithms.
  • We provide a comparative analysis of the performance of five MMD algorithms, each based on a different MAPF algorithm, shedding light on their applicability to coordinating robots leveraging diffusion models for planning.
  • We show that we can scale our approach to arbitrarily large and diverse maps by learning and composing multiple diffusion models for each robot.
Our experimental results from varied motion planning problems in simulated scenarios motivated by logistics environments suggest that our approach scales favorably with both the number of agents and the size of the environment when compared to alternatives.
@@ 2.3 Planning with Diffusion Models
$\operatorname*{arg\,max}_{\boldsymbol{\tau}^{i}}\log p(\boldsymbol{\tau}^{i}|\mathcal{O})=\operatorname*{arg\,max}_{\boldsymbol{\tau}^{i}}\left(\log p(\mathcal{O}|\boldsymbol{\tau}^{i})+\log p(\boldsymbol{\tau}^{i})\right)$ (1)
$=\operatorname*{arg\,min}_{\boldsymbol{\tau}^{i}}\left(-\log p(\mathcal{O}|\boldsymbol{\tau}^{i})-\log p(\boldsymbol{\tau}^{i})\right)=\operatorname*{arg\,min}_{\boldsymbol{\tau}^{i}}\left(\mathcal{J}(\boldsymbol{\tau}^{i})-\log p(\boldsymbol{\tau}^{i})\right)$
@@ 3.2 Constraining Strategies
To determine when and where to apply constraints on diffusion models, MMD draws on MAPF strategies like CBS and PP. We propose five MMD variants, each inspired by a state-of-the-art search algorithm. Alg. 1 provides a summary of these methods and we elaborate upon them here.
@@ 3.3 Sequencing Diffusion Models for Long Horizon Planning
Diffusion models have shown notable success in learning trajectory distributions within specific contexts. However, they face challenges in modeling complex trajectory distributions and generalizing to diverse contexts (e.g., significantly different obstacle layouts). To tackle this issue, we propose utilizing an ensemble of local diffusion models rather than a single model for a single robot. Each local model is trained to capture the trajectory distribution of a particular motion pattern within a local map segment, and we combine multiple models sequentially for long-horizon single-robot planning. This approach offers three advantages: (1) Easier learning with simple, demonstrable motion patterns. (2) Better generalization across different contexts. (3) Effective scaling for large maps, aiding long-term planning.
@@ 4 Experimental Analysis
Figure 3: Analysis of success rates and data adherence scores in randomly generated planning queries. We report on all five MMD instantiations, each influenced by a constraint-based search algorithm, and a MAPF method with and without a learned cost map. The left column shows our test maps, the middle column compares success rates across $10$ trials per robot count, and the right column presents the average data adherence scores. Sec. 4 discusses the experiments and results.
@@
4.1 Decoupling Scales Multi-Robot Diffusion Planning
4.2 MMD Outperforms MAPF with Learned Cost Maps
From a search-based planning viewpoint, a compelling way to integrate motion data into multi-agent planning is by reducing MRMP to MAPF and forming cost maps to direct algorithms like ECBS (Cohen et al., 2016). This sacrifices dynamic feasibility, limiting solutions to a fixed graph, but can still offer a desirable homotopy class and be assessed for data compliance. Our second experiment set evaluates this method’s potential to produce data-compliant motions (see Fig. 3(j)). We created 180 motion planning problems, 10 per group size across three maps, and assessed two search-based methods: A*Data-ECBS and A*-ECBS. The former plans single-robot paths with A* using map-specific statistical cost maps, where edges have lower costs if those are visited more frequently by the map’s dataset trajectories (discretized to map cells). The latter uses uniform costs (i.e., has no knowledge of the data) and is reported only to provide context for A*Data-ECBS’s performance.
We find that integrating motion data into statistical cost maps yields mixed outcomes: it enhances adherence in the Highways map by around $35\%$ compared to A*-ECBS but decreases success rates. In other maps, this approach finds collision-free paths, but falls short in managing complex motion distributions. In contrast, our results for running MMD methods for the same planning problems show that those can solve complex trajectory generation tasks with high data adherence scores across all methods, and for some, with high success rates too.
@@ 4.2 MMD Outperforms MAPF with Learned Cost Maps
Figure 4: Scalability tests for our five MMD methods in free-space MRMP. Starts and goals were chosen to induce robot collisions. The top row shows the circle setup, where robots swap positions between opposite points on the perimeter. The weave setup results are below, with robots exchanging positions along uniformly spaced boundary points. Length is zero for failed problems (in MMD-PP, caused by yielding invalid solutions, and in the rest, by computing for longer than 240 seconds).
@@
Analyzing MMD Constraining Strategies. Further analysis reveals a trend familiar from the MAPF literature. MMD-CBS struggles with scalability, reaching the time limit with more than $9$ robots. MMD-ECBS significantly outperforms MMD-CBS, and accelerated versions further improve performance. MMD-PP finds efficiency in requiring only one inference pass per robot, however, because the constraints on diffusion models are soft, trajectory generation queries are not guaranteed to completely avoid higher-priority robots, and as such MMD-PP may fail to produce collision-free solutions. This is reflected in lower success rates in congested maps. In contrast, CBS-based MMD methods only failed by exceeding our 60-second runtime limit. MMD-xECBS outperformd other MMD algorithms in success rates and matched MMD-ECBS in data adherence.
To test MMD’s MRMP solving ability without regard to data adherence, we created two free-space experiments focusing on robot interactions (Fig. 4(f)). In the circle setup, robots move between opposite points on a circle, likely colliding at the center. In the weave setup (inspired by Tajbakhsh et al. (2024)), robots begin on opposite points of a square, aiming to switch places. CBS-based methods were challenged in circle since incrementally constraining the center region is time consuming. MMD-PP’s stronger constraints navigated around congestion and solved more problems, however, occasionally failed to produce valid solutions. In weave, where navigating around congestion is more difficult, CBS-based methods generated collision-free trajectories more effectively.
@@
4.3 Sequenced Diffusion Models for Long-Horizon Planning
@@ 5 Related Work
Multi-Robot Motion Planning. Due to the PSPACE-hardness of MRMP (Hopcroft & Wilfong, 1986), most practical algorithms make various approximations to relax the planning problem. Some of the most successful MRMP methods reduce the problem to a multi-agent path finding problem (MAPF) (Stern et al., 2019) by discretizing state and time and assuming zero-volume shapes, constant velocities and rectilinear movements (Chen et al., 2021). However, this introduces a gap between the MAPF formulation and MRMP which is problematic in many domains. To relax the restriction of planning on a grid, recent works (Solis et al., 2021; Hönig et al., 2018) have generalized MAPF algorithms to continuous state spaces by using sampling-based motion planners to construct probabilistic roadmaps (PRM)  (Kavraki et al., 1996) for robots.
@@
More generally, coupled MRMP algorithms, like Sanchez & Latombe (2002); Solovey et al. (2016) apply sampling-based planners like PRM and RRT (Karaman & Frazzoli, 2011) to the composite configuration space of all the robots. This approach guarantees completeness but struggles to scale with the number of robots as the configuration space grows exponentially. Conversely, decoupled algorithms like prioritized planning (Erdmann & Lozano-Perez, 1987) and Leroy et al. (1999) generate robot motions one after another, fixing each plan and regarding them as dynamic obstacles in subsequent planning iterations. This allows them to quickly find solutions for large numbers of robots, but sometimes fail to find any solution even if one exists, i.e., they are often incomplete. Seeking to balance theoretical guarantees with practical efficiency, hybrid algorithms like CBS (Sharon et al., 2015) plan independently for every robot and loosely couple them to resolve conflicts.
@@
Planning with Diffusion Models. Lately, there has been a surge of interest in applying diffusion models (Sohl-Dickstein et al., 2015) to solve sequential decision making tasks, including planning and reinforcement learning (Ubukata et al., 2024). Diffuser (Janner et al., 2022) first proposed the idea of using diffusion models for trajectory planning and showed how classifier-guided sampling and image inpainting can be used for adaptation at test time. Recent works have used diffusion models as priors for single-robot motion planning (Carvalho et al., 2023), for learning visuomotor policies (Chi et al., 2024), and for offline decision-making (Ajay et al., 2023). Most of these works have been limited to single-agent planning with two notable exceptions.  Jiang et al. (2023) learn a joint motion distribution for multi-agent motion prediction using a diffusion model. By contrast, we focus on generating kinodynamically feasible and collision-free trajectories without having to learn a joint distribution. Concurrent work by Mishra et al. (2024) uses spatial-temporal factor graphs to compose modular generative models for solving long-horizon bimanual coordination tasks. By contrast, our approach can scale diffusion models to dozens of robots and complex environments by leveraging multi-agent search algorithms.
@@ 6 Conclusion
In this paper, we present Multi-robot Multi-model planning Diffusion (MMD), a multi-robot motion planner that learns to generate smooth collision-free trajectories for dozens of robots in complex environments. Our key contribution is showing how single robot diffusion models can be effectively combined with state-of-the-art multi-agent search algorithms to generate data-driven multi-robot trajectories. By learning only single-robot diffusion models, MMD simplifies data requirements and breaks the curse of dimensionality plaguing approaches that learn from multi-robot data. Additionally, by learning generative models of robot trajectories, MMD overcomes many of the limitations of popular model-based MAPF algorithms, such as state discretization, known cost function, and constant velocities. We believe MMD opens up exciting avenues for future work on combining the strengths of search algorithms and diffusion planning.
Limitations. We are excited about the potential of MMD to push forward multi-robot coordination and collaboration and offer a few avenues for future work. First is allowing for more flexibility in MMD’s planning horizon. As it stands, MMD generates trajectories of fixed duration—a limitation it inherits from diffusion planning. Sequencing models helps extend the time horizon but still, it must be determined a priori. Second, we believe that the CBS-based MMD methods can be greatly accelerated, mostly through parallelization of the high-level search. Finally, we believe the frontier of MRMP lies in collaboration. Currently, MMD focuses on coordinating robots, seeking to produce collision-free data-driven trajectories. Carrying out collaborative tasks is an interesting next step.
@@ Acknowledgments
The work was supported by the National Science Foundation under Grant IIS-2328671 and the CMU Manufacturing Futures Institute, made possible by the Richard King Mellon Foundation. We thank Mohit Sharma and Ilan Mitnikov for fruitful discussions.

### v2

@@ Abstract
Diffusion models have recently been successfully applied to a wide range of robotics applications for learning complex multi-modal behaviors from data. However, prior works have mostly been confined to single-robot and small-scale environments due to the high sample complexity of learning multi-robot diffusion models. In this paper, we propose a method for generating collision-free multi-robot trajectories that conform to underlying data distributions while using only single-robot data. Our algorithm, Multi-robot Multi-model planning Diffusion (MMD), does so by combining learned diffusion models with classical search-based techniques—generating data-driven motions under collision constraints. Scaling further, we show how to compose multiple diffusion models to plan in large environments where a single diffusion model fails to generalize well. We demonstrate the effectiveness of our approach in planning for dozens of robots in a variety of simulated scenarios motivated by logistics environments.
@@ 1 Introduction
In this paper, we propose a data-efficient and scalable multi-robot diffusion planning algorithm, Multi-robot Multi-model planning Diffusion (MMD), that addresses both these challenges by combining constraint-based MAPF planners with diffusion models. Importantly, our approach calls for learning only single-robot diffusion models, which does away with the difficulty of obtaining multi-robot interaction data and breaks the curse of dimensionality. MMD generates collision-free trajectories by constraining single-robot diffusion models using our novel spatio-temporal guiding functions and choosing constraint placement via strategies inspired by MAPF algorithms. Our contributions in this paper are threefold: (1) We propose a novel data-efficient framework for multi-robot diffusion planning inspired by constraint-based search algorithms. (2) We provide a comparative analysis of the performance of five MMD algorithms, each based on a different MAPF algorithm, shedding light on their applicability to coordinating robots leveraging diffusion models for planning. (3) We show that we can scale our approach to arbitrarily large and diverse maps by learning and composing multiple diffusion models for each robot. Our experimental results from varied motion planning problems in simulated scenarios motivated by logistics environments suggest that our approach scales favorably with both the number of agents and the size of the environment when compared to alternatives. Video demonstrations and code are available at https://multi-robot-diffusion.github.io/.
@@ 2.3 Planning with Diffusion Models
$\operatorname*{arg\,max}_{\boldsymbol{\tau}^{i}}\log p(\boldsymbol{\tau}^{i}|\mathcal{O})=\operatorname*{arg\,min}_{\boldsymbol{\tau}^{i}}\left(\mathcal{J}(\boldsymbol{\tau}^{i})-\log p(\boldsymbol{\tau}^{i})\right)$ (1)
@@ 3.2 Constraining Strategies
To determine when and where to apply constraints on diffusion models, MMD draws on MAPF strategies like CBS and PP. We propose five MMD variants, each inspired by a state-of-the-art search algorithm. Alg. 1 provides a summary of these methods and we elaborate upon them here 222 In MMD, we use the search or prioritization logic found in MAPF algorithms for placing “strong” constraints on diffusion models, while all other aspects of MMD are more loosely inspired by MAPF algorithms. .
@@ 3.3 Sequencing Diffusion Models for Long Horizon Planning
Diffusion models have shown notable success in learning trajectory distributions within specific contexts. However, they face challenges in modeling complex trajectory distributions and generalizing to diverse contexts (e.g., significantly different obstacle layouts). We propose utilizing an ensemble of local diffusion models for each robot to facilitate varying-context planning. Each local model is trained to capture a particular motion pattern, i.e., a trajectory distribution generated by a hidden cost function defined by a specific task dataset. For example, near a conveyor belt, we can define a motion pattern requiring robots to pass through either the top corridor right-to-left, or the bottom corridor left-to-right. By sequentially combining multiple local models, each corresponding to a local map segment, we enable long-horizon single-robot planning that is easier to learn, generalizes well to different contexts, and scales effectively to large maps.
@@ 4 Experimental Analysis
4.1 Decoupling Scales Multi-Robot Diffusion Planning
4.2 MMD Outperforms MAPF with Learned Cost Maps
From a search-based planning viewpoint, a compelling way to integrate motion data into multi-robot planning is by reducing MRMP to MAPF and forming cost maps to direct algorithms like ECBS (Cohen et al., 2016). This sacrifices dynamic feasibility, limiting solutions to a fixed graph, but can still offer a desirable homotopy class and be assessed for data compliance. Our second experiment set evaluates this method’s potential to produce data-compliant motions (see Fig. 3(j)). We created 180 motion planning problems, 10 per group size across three maps, and assessed two search-based methods: A*Data-ECBS and A*-ECBS. The former plans single-robot paths using A* (Nilsson, 1982) with statistical cost maps, where edge costs are lower if map dataset trajectories frequently visit those areas. The latter uses uniform costs (i.e., has no knowledge of the data) and is reported only to provide context for A*Data-ECBS’s performance.
Integrating motion data into statistical cost maps shows mixed results: it improves Highways map performance by $35\%$ over A*-ECBS but reduces success rates. In other maps, it finds collision-free paths but struggles with complex motion distributions. MMD methods, by contrast, effectively generate trajectories with high data adherence and often high success rates. As shown in A.2.1 and Table 2, planning time and solution quality correlate: when A*Data-ECBS matches MMD’s data adherence, their planning times align, but faster A*Data-ECBS solutions come at the cost of lower quality. We believe there is significant potential to improve the computational efficiency of MMD through parallelization. However, we have left such optimizations for future work.
@@ 4.2 MMD Outperforms MAPF with Learned Cost Maps
Figure 3: Analysis of success rates and data adherence scores, in randomly generated planning queries, of all MMD instantiations and a MAPF method with and without a learned cost map. The left column shows our test maps, the middle column compares success rates across $10$ trials per robot count, and the right column presents the average data adherence scores.
@@
4.2.1 Analyzing MMD Constraining Strategies.
Further analysis reveals a trend familiar from the MAPF literature. MMD-CBS struggles with scalability, MMD-ECBS significantly outperforms it, and accelerated versions further improve performance. MMD-PP finds efficiency in requiring only one inference pass per robot, however, because the constraints on diffusion models are soft, trajectory generation queries are not guaranteed to completely avoid higher-priority robots, and as such MMD-PP may fail to produce collision-free solutions. This is reflected in lower success rates in congested maps. In contrast, CBS-based MMD methods only failed by exceeding our 60-second runtime limit. MMD-xECBS outperformd other MMD algorithms in success rates and matched MMD-ECBS in data adherence.
To test MMD’s MRMP solving ability without regard to data adherence, we created two free-space experiments focusing on robot interactions (Fig. 4(f)). In the circle setup, robots move between opposite points on a circle, likely colliding at the center. In the weave setup (inspired by Tajbakhsh et al. (2024)), robots begin on opposite points of a square, aiming to switch places. CBS-based methods were challenged in circle since incrementally constraining the center region is time-consuming. MMD-PP’s stronger constraints navigated around congestion and solved more problems, however, occasionally failed to produce valid solutions. In weave, where navigating around congestion is more difficult, CBS-based methods generated collision-free trajectories more effectively.
4.3 Sequenced Diffusion Models for Long-Horizon Planning
@@
Figure 4: Scalability tests in high-congestion free-space MRMP. Circle (top row) asks robots to swap positions between opposite points on the perimeter. Weave (below), asks robots to exchange positions along uniformly spaced boundary points. Length is zero for failed problems (MMD-PP failures were due to yielding invalid solutions, and other methods failed by exceeding 240 seconds).
@@ 5 Related Work
Multi-Robot Motion Planning. Many MRMP algorithms (Sanchez & Latombe, 2002; Solovey et al., 2016) approach the problem by treating it as a coupled system – applying sampling-based planners such as PRM (Kavraki et al., 1996) and RRT (Karaman & Frazzoli, 2011) to the composite configuration space of all robots. While this approach guarantees probabilistic completeness, it struggles to scale with the number of robots due to the exponential growth of the configuration space. Given the PSPACE-hardness of MRMP (Hopcroft & Wilfong, 1986), many practical algorithms introduce approximations to make the planning problem more tractable. Some of the most successful MRMP methods transform the problem into a multi-agent path finding (MAPF) problem (Stern et al., 2019) through state and time discretization  (Li et al., 2019; Hönig et al., 2018).
@@
Decoupled algorithms like prioritized planning (Erdmann & Lozano-Perez, 1987) and Leroy et al. (1999) generate robot motions one after another, fixing each plan and regarding them as dynamic obstacles in subsequent planning iterations. This allows them to quickly find solutions for large numbers of robots, but sometimes fail to find any solution even if one exists, i.e., they are often incomplete. Seeking to balance theoretical guarantees with practical efficiency, hybrid algorithms like CBS (Sharon et al., 2015) and its variants (Barer et al., 2014; Li et al., 2021) decouple the MRMP into two levels – high-level, in which they resolve conflicts between robots, and low-level, in which they plan motion for each robot independently. A* and its derivatives are the de facto standard for motion planning for navigation, serving as the low-level search algorithm in many of these state-of-the-art approaches, as they are efficient, complete, and bounded sub-optimal.
@@
Planning with Diffusion Models. Lately, there has been a surge of interest in applying diffusion models (Sohl-Dickstein et al., 2015) to solve sequential decision making tasks, including planning and reinforcement learning (Ubukata et al., 2024). Diffuser (Janner et al., 2022) first proposed the idea of using diffusion models for trajectory planning and showed how classifier-guided sampling and image inpainting can be used for adaptation at test time. Recent works have used diffusion models as priors for single-robot motion planning (Carvalho et al., 2023), for learning visuomotor policies (Chi et al., 2024), and for offline decision-making (Ajay et al., 2023). Most of these have been limited to single-agent planning with two notable exceptions.  Jiang et al. (2023) learns a joint motion distribution for multi-agent motion prediction and concurrent work by Mishra et al. (2024) uses spatial-temporal factor graphs to compose modular generative models for solving long-horizon bimanual tasks. By contrast, we focus on generating kinodynamically feasible and collision-free trajectories for dozens of robots and complex environments without learning a joint distribution.
@@ 6 Conclusion
In this paper, we present MMD, a multi-robot motion planner that learns to generate smooth collision-free trajectories for dozens of robots in complex environments. Our key contribution is showing how single robot diffusion models can be effectively combined with search algorithms to generate data-driven multi-robot trajectories. By learning only single-robot diffusion models, MMD simplifies data requirements and breaks the curse of dimensionality plaguing approaches that learn from multi-robot data. Additionally, by learning generative models of robot trajectories, MMD overcomes many of the limitations of popular model-based MAPF algorithms, such as state discretization, known cost function, and constant velocities. We believe MMD opens up exciting avenues for future work on combining the strengths of search algorithms and diffusion planning.
Limitations. We are excited about the potential of MMD to push forward multi-robot coordination and collaboration and offer a few avenues for future work. First is combining MMD with decentralized and windowed multi-robot planning algorithms, like Gaussian Belief Propagation (GBP) (Patwardhan et al., 2023), whose collision avoidance signals could be incorporated into single-robot diffusion guidance functions. We discuss this research direction in Sec. A.1.1. Second, we believe that the CBS-based MMD methods can be greatly accelerated, mostly through parallelization of the high-level search. Finally, we believe the frontier of MRMP lies in collaboration. Currently, MMD focuses on coordinating robots, seeking to produce collision-free data-driven trajectories. Carrying out collaborative tasks is an interesting next step.
@@ Acknowledgments
We thank Mohit Sharma and Ilan Mitnikov for fruitful discussions. This work was supported by the National Science Foundation under grant IIS-2328671 and grant #1955361, as well as by the Office of Naval Research (ONR) under REPRISM MURI N000142412603 and ONR grant #N00014-22-1-2592. Partial funding was also provided by a gift from Amazon and the Robotics and AI Institute.
@@ A.1 Additional Algorithmic Discussion
A.1.1 Beyond Full-Horizon Planning
In this work, we focus on scenarios where it is possible to carry out centralized planning. That is, a single algorithm generates full-horizon trajectories for all robots, and robots later execute these trajectories. This formulation is common in robotics and the planning community more broadly, however, it has some limitations. For example, this setup requires that the motion of all moving obstacles be known a priori, that all robots be able to carry out their trajectories as prescribed, and that time for offline planning be available. In the real world, such information or resources may not always be available (e.g., when robots operate next to humans). In this section, we provide a brief introduction to another class of multi-robot planning algorithms—decentralized and windowed algorithms—that address some of these challenges. We also explore an exciting avenue for future work: combining these algorithms with MMD to mitigate the limitations of each approach.
The structure of decentralized and windowed multi-robot planning algorithms differs fundamentally from full-horizon planning algorithms in that they break the “plan-then-act” paradigm (Patwardhan et al., 2023; Van Den Berg et al., 2011; Okumura et al., 2019). While full-horizon planners first generate a set of trajectories for all robots and then robots execute them as prescribed, windowed algorithms instead ask each robot to plan a short trajectory for itself, execute it, observe the new state of the world, and repeat. This setup gives up on global optimality in favor of allowing faster planning cycles and decentralized computation. In that, it removes the reliance on a central planner that implicitly also assumes perfect communication between robots. Instead, planned trajectories are communicated and negotiated between neighboring robots. We believe that MMD could benefit from this structure, as it enables adaptation to dynamic environments and reduces the combinatorial complexity of MAPF algorithms by which our proposed MMD algorithms are inspired. However, the best approach for adapting MMD to decentralized and windowed settings remains unclear.
One algorithm that could shed light on how to step towards decentralized and windowed data-driven trajectory generation with MMD is Gaussian Belief Propagation Planner (GBP) (Patwardhan et al., 2023). This planner is a recently proposed decentralized and windowed algorithm for multi-robot planning that frames short-horizon planning as inference. A particular similarity between GBP and MMD is the way single-robot plans influence each other to achieve collision-free multi-robot plans. In both MMD and GBP, soft constraints are imposed on single-robot time-discretized trajectories (either full-horizon or a short-horizon window) to guide their trajectory generation processes towards favorable regions in trajectory space (e.g., collision-free, respecting dynamics, etc.). Given this similarity, it is reasonable to believe that by incorporating signals from GBP’s factor graph into MMD’s guidance functions, MMD could be adapted to the decentralized and windowed setting.